find the real solution of each equation. where necessary, round to the nearest hundredth.
1.)x^2+3x-4=0
2.)x^2+2x-3=0
3.)x^3+2x^2+1x=0

Answer by Alan3354(69443)

find the real solution of each equation. where necessary, round to the nearest hundredth.
1.)x^2+3x-4=0
2.)x^2+2x-3=0
3.)x^3+2x^2+1x=0
-------------------
All 3 can be factored. That's the simplest method.